OpenStudy (anonymous):

Line segment EF has endpoints at E(-2, 1) and F(-5, 5). Line segment GH has endpoints at G(-1, -2) and H(3, 1). What are their lengths?

OpenStudy (anonymous):

You can find a right triangle made by each line segment. For example, EF has a rise (y-values) of (5-1) = 4 and a run (x-values) of (-5-(-2)) = -3. So square both the numbers and square root the answer. So 4^2+(-3)^2= 25. Square root of 25 = 5.

OpenStudy (anonymous):

Follow the same process for GH also.

OpenStudy (anonymous):

oh ok i gotcha thanks! : )